OpenStudy (anonymous):

A pair of equations is shown below: y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution.

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

Do you know what elimination and substitution method are?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Substitution is when you replace one variable from the 1st equation using the 2nd equation. Elimination is when you manipulate the two equations, so that when you add those two , one variable will be eliminated. So looking at your equations, substitution is the easiest method. \(\sf \color{red}{y} = 8x − 9\) \(\sf \color{red}{y = 4x − 1}\) replace y in the 1st equation with the 2nd equation it will be like this: \(\sf \color{red}{4x − 1}= 8x − 9\) dis this part clear?

OpenStudy (anonymous):

oh ok yea now i know and then sovle right

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

is that right @Data_LG2

OpenStudy (anonymous):

let's see 4x−1=8x−9 9-1=8x-4x 8=4x hmm I don't think so..

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

is that right @Data_LG2

OpenStudy (anonymous):

yes. do you understand it though?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

can you help me find y

OpenStudy (anonymous):

to find y, just plug the x value that you found in any of the two equations given

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

@Data_LG2 can you help me

OpenStudy (anonymous):

choose one of the equations given

OpenStudy (anonymous):

tell me which one you choose 1st or 2nd equation?

OpenStudy (anonymous):

1st

OpenStudy (anonymous):

can you help me???

OpenStudy (anonymous):

alright. sorry i was afk y=8x−9 x = 2 from your calculation. replace x in the 1st equation by 2 y= 8(2) - 9 then evaluate, what will you get?

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

is that right @Data_LG2

OpenStudy (anonymous):

what is 8 x 2?

OpenStudy (anonymous):

then subtract 9 from the product of those two, what will you get?